OpenStudy (anonymous):

Find the solutions to 0 ≥ x4 – 1

OpenStudy (maheshmeghwal9):

is this ur question? \[0 \ge x^4-1.\]???????????????????????????????????????????????????????????????????????????

OpenStudy (maheshmeghwal9):

@chase3 ????????????????????????????????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so am i suppose to start my x values with -1, 0, 1 ,2

OpenStudy (anonymous):

so im only going to have two x values

OpenStudy (maheshmeghwal9):

no a set of values

OpenStudy (maheshmeghwal9):

i think u need more clarification.

OpenStudy (anonymous):

yea that is what im talking about ... explain to me how i would put that on a function chart

OpenStudy (maheshmeghwal9):

wait a minute i m going wrong

OpenStudy (maheshmeghwal9):

We would have 4 values -1, +1, -i, +i . Ok!

OpenStudy (anonymous):

ok

OpenStudy (maheshmeghwal9):

I think this question is going on high level than mine! I don't know wt to do of imaginary critical points that is - i & +i. So i think @waterineyes Plz help:D

OpenStudy (maheshmeghwal9):

We have done only this much \[x^4-1 \le 0.\]\[\implies (x^2+1)(x^2-1) \le 0.\]\[\implies (x^2+1)(x+1)(x-1) \le 0.\]So critical point are : -\[+1, -1, +i , -i.\] But wt to do now????????????

OpenStudy (anonymous):

i think your'e just suppose to plug in the x values into the equation

OpenStudy (maheshmeghwal9):

i think strictly NO

OpenStudy (maheshmeghwal9):

we want a representation of this inequality to get the correct solution set.

OpenStudy (maheshmeghwal9):

but how to represent it?

OpenStudy (maheshmeghwal9):

@Ishaan94 @amistre64 @mathslover @phi Plz help:)

OpenStudy (maheshmeghwal9):

@waterineyes too plz :)

OpenStudy (anonymous):

i really do think we're over thinking the question

mathslover (mathslover):

\[\large{0\ge x^4-1}\] \[\large{0+1 \ge x^4-1+1}\] \[\large{1 \ge x^4}\] \[\large{1\ge (x)^4}\] \[\large{\sqrt[4]{1}\ge x}\] \[\large{1\ge x}\]

mathslover (mathslover):

thanks for tagging me here @maheshmeghwal9

OpenStudy (anonymous):

but im pretty sure the answer is suppose to be in a table chart

OpenStudy (phi):

If you are allowing complex values, the solution is the complex numbers less than or equal to the unit circle.

mathslover (mathslover):

either x = 1 or x<1

OpenStudy (phi):

If you are only dealing with real numbers then -1≤ x ≤ 1

OpenStudy (anonymous):

so @phi you think that is the answer

OpenStudy (phi):

try it. (-1)^4 - 1 = 0 which agrees with ≤ 0 clearly fractions less than 1 raised to the 4th power are < 1, and we match the condition

OpenStudy (phi):

if you allow complex numbers then x= A exp(i theta) where |A|≤ 1, 0≤theta≤ 2pi will meet the condition.

OpenStudy (anonymous):

alright will do